In a recent electin Corrine Brown received 13,356 more votes than Bill Randll. If thetotal number of votes was 119,626, find the number of votes for each candidate.
Corrine Brown received_________ votes?
Bill Randall received _____ votes?

Answer by mananth(16946)

In a recent electin Corrine Brown received 13,356 more votes than Bill Randll. If thetotal number of votes was 119,626, find the number of votes for each candidate.
bill - x votes
corrine = x+13356
x+x+13356= 119626
2x=119626-13356
2x=106270
x=53,135
Corrine Brown received__66,491_______ votes?
Bill Randall received __53,135___ votes?
...
m.ananth@hotmail.ca